OpenStudy (anonymous):

I give medals---Solve: 7/2t + 2 = 4 + 7t A. t = -7/4 B. t = -4/7 c. All numbers D. No solution

OpenStudy (anonymous):

B. t= -4/7 stated both sides equal 0. what does that mean?

OpenStudy (anonymous):

is b the answer?

OpenStudy (anonymous):

what does no solution mean

OpenStudy (cwrw238):

the question is a bit ambiguous is the left hand side (7/2t ) + 2 or 7 / (2t + 2) ?

OpenStudy (cwrw238):

or of course it could be (7/2) t or 3.5t

OpenStudy (cwrw238):

i think the last one is most likely

OpenStudy (cwrw238):

in which case it simplifies to 3.5t = -2

OpenStudy (cwrw238):

and t = -4/7 is the correct answer

OpenStudy (texaschic101):

(7/2)t + 2 = 4 + 7t --- multiply everything by 2 7t + 4 = 8 + 14t 7t - 14t = 8 - 4 -7t = 4 t = -4/7

OpenStudy (cwrw238):

to avoid confusion the equation should be written as texaschic wrote it